1.ABC Company currently has a cash cycle of 238 days. The company went through a management change. The new management has changed the working capital policies for the company. The new policies will decrease the inventory period by 24 days, decrease the accounts receivable period by 30 days, and decrease the accounts payable period by 45 days. Calculate the number of days in the new cash cycle after the above changes become effective? That is, what is the new cash cycle?

2.The terms of the sale were 5/11, net 61. What is the effective annual rate of interest? Convert your answer to percentage and round off to two decimal points. Do not enter % in the answer box.

3.ABC Company sells 32,212 chairs a year. The carrying cost per unit is $2.66 and the fixed costs per order are 198.34 per order. The company orders 358 units each time. The chairs are sold out before they are restocked. What is the economic order quantity?Enter your answer rounded off to two decimal points. Do not enter comma in the answer box.
Answer #1

1.New Cash cycle Days= 238-45-24+30 =199 days

In this case we are paying 5% more for delayig the payment by 61-11 - so days 5 x 100 = 5.26% effective rate. - for so days APlease leave a like if the answer helped you. Thank u!